OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. x/2+4/5=9/12

OpenStudy (anonymous):

/5: 5 , 10, 15, 20, 25, 30 9/12=3/4, and x/2=2x/4 so /4: 4, 8, 12, 16, 20, 24, 28, 32

OpenStudy (dumbcow):

hmm in general you multiply all the prime factors while ignoring repeats --> 2* 5 * (2*2*3) notice 2 and 12 share factor of 2 so do not include the first "2" 5*(2*2*3) = 60 LCD = 60

OpenStudy (anonymous):

The perimeter of a rectangular garden is 52 feet. The length of the garden is 4 feet less than twice the width. What is the width of the garden?